OpenStudy (anonymous):

A wire in the shape of a circle encloses the area 38.5 sqcm. if the wire is rebent to form a square,find the of the square.

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Find the ? of the square...

OpenStudy (anonymous):

area

terenzreignz (terenzreignz):

Okay, let's go for basics... If you have an 8cm wire, bent into a square, how long would one side be?

OpenStudy (anonymous):

first i have to put the formula for area of circle pir^2= 38.5 from here i got value of r = 7/2

terenzreignz (terenzreignz):

You got an exact fraction? I find this hard to believe...

OpenStudy (anonymous):

then i used circumference of a circle = 2 *pi *r=2*pi*(7/2)=22cm now, circumference of a cirlce = perimeter of a wire=22cm

OpenStudy (anonymous):

plse tell till now i have solved the sum correct

terenzreignz (terenzreignz):

Approximately, yeah... Are you sure you needed help with this? :D

OpenStudy (anonymous):

yes.. i m not sure that my answer is correct or not

terenzreignz (terenzreignz):

Well, I'll vouch for you, on the grounds that you understand that that's just an approximation (as if you can get an exact answer with pi involved, but whatever...)

OpenStudy (anonymous):

i got the area of a square = 121/4 cm^2

terenzreignz (terenzreignz):

That seems about right; well done :)

OpenStudy (anonymous):

thank you